# Data Mixing Laws: Optimizing Data Mixtures by Predicting Language Modeling Performance

### v1

@@ 1 Introduction
Pretraining data for large language models (LLMs) are typically a mixture of multiple domains, varying from English to minority languages (Doddapaneni et al., 2021; Li et al., 2023), from casual dialogs to formal academic writings (Taylor et al., 2022), and from texts to modalities like images and speeches (Zhan et al., 2024), among others. These data interplay with each other, showing complex interchangeable, unrelated, or contradictory relationships (Guo et al., 2024). This necessitates adjusting the mixture proportions of training data to balance the model capabilities while harnessing synergies across domains, thus enhancing the competence of the outcome models, as highlighted by extensive practical experience (Gao et al., 2020; Gururangan et al., 2020; Zhou et al., 2023; Xie et al., 2024a).
Nonetheless, it remains elusive to figure out an ideal training data mixture. Most existing practices tune the mixture through heuristics to upsample a proportion of high-quality or underrepresented data without disclosing the concrete criteria in detail (Gao et al., 2020; Touvron et al., 2023a; Bai et al., 2023; Bi et al., 2024) and it is hard to predate whether these data strategies are effective before finishing the training run. Encouraged by advances in scaling laws that show model losses on a given set of evaluation data are quantitatively predictable for a wide range of variables (Kaplan et al., 2020; Hoffmann et al., 2022), we wonder whether this also holds for mixture proportions, so that we can estimate the outcome model performance given any mixture before actually training on them, including the desired one that reaches minimum loss.
In this paper, we answer this proposition affirmatively. We find that, given a mixture of $M$ domains, an exponential function over the linear combination of the proportions, i.e.,
@@ 2 Background
We briefly review the pretraining process of large language models and summarize key findings from neural scaling laws, then we formalize the problem we study. For more related work, please refer to Sec. 6.
@@
For a wide range of factors $x$, scaling laws (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022) show that their effect on the loss $L$ of a pretrained model follows power laws
@@
where $c$, $k$, and $\alpha$ are parameters to fit and $x$ can be model sizes, numbers of training data, training steps222 The training step law (Kaplan et al., 2020) refers to the function between the intermediate training step and its corresponding loss in one single training run. For clarity, we highlight its difference from the data scaling law that predicts the early exit loss (Kaplan et al., 2020) or the loss obtained with optimal hyperparameters for each data budget (Hoffmann et al., 2022). , and the amount of computation. Previous experience (Alabdulmohsin et al., 2022; OpenAI, 2023; Bi et al., 2024; Su et al., 2024) highlights the impressive predictability of scaling laws. Specifically, Eqn. 3 fitted on a collection of small models, training data, or computation can extrapolate to precisely predict the test loss of larger cases over orders of magnitudes. This enables practitioners to estimate the performance of a pretrained large language model without actually finishing the expensive runs. Recent development further shows various functional relationships between the performance of language models and a broader range of factors, including transfer learning (Hernandez et al., 2021), sparse architectures (Frantar et al., 2023), and repeated data (Muennighoff et al., 2024), consolidating the predictability of language model performance.
@@ 3.1 Pilot Study on Domain Losses under Two-domain Mixtures
Results in Fig. 2 reveal the quantitative predictability of domain losses given the domain proportions. We encouragingly find that, for checkpoints with the same size and trained with the same number of steps, after subtracting a shared constant [^3] The loss function of language models is lower bounded by the entropy of the evaluation data theoretically (Bishop, 2006; Henighan et al., 2020). Thus, it is common to have constant terms in scaling laws, known as irreducible loss. , their domain losses in the log scale demonstrate a linear relationship to the domain proportion. This holds for both domains in our experiments. The result indicates that with other factors fixed, the domain losses of a pretrained language model regarding the domain proportion precisely fit into an exponential law
@@ 3.2 Extension to Domain Losses Trained on Multi-domain Mixtures
We train on the mixtures of GitHub, Pile-CC, and Books3 subset from the Pile for a total of 30B tokens and evaluate the model on the three domains, respectively. For specific mixtures, we grid search from $\{0,0.125,0.25,\dots,0.875,1\}^{3}$ and retain valid ones in which three proportions sum up to 1 and do not use up all tokens in any of the domains [^5] The GitHub and Books3 subset in the Pile do not contain as many as 30B tokens after deduplication., which results in 32 mixtures in total.
@@
  • Symmetry. Any exchanging of variables should not change the functional form.
The second principle stems from the intuition to avoid introducing any domain-specific bias. Together, the two principles lead to candidate functions that replicate the exponential term in Eqn. 6 for each training domain and combine them through operations that adhere to commutative law.
@@
To provide more intuition, we discuss the meanings of the coefficients in Eqn. 7. In general, $k_{i}>0$, thus the exponential term is always positive and the prediction loss is strictly greater than the constant $c$. Hereby, $c_{i}$ represents losses that are not reducible by adjusting the data mixture. $t_{ij}$, depending on both training domain $i$ and validation domain $j$, shows the interaction between them. A negative $t_{ij}$ indicates that training data of domain $j$ helps reduce validation loss on domain $i$ and vice versa. Besides, rewriting the coefficient $k_{i}$ into $\exp(\log k_{i})$ makes it a bias term on the exponent which adjusts the aggregate effect of all training domains.
3.3 Predicting Losses of Any Validation Mixture
@@
A remaining limitation is that we still need to acquire the components of validation data $s_{1\dots K}$ in advance. This can be inconvenient if the validation set is collected separately from the training ones. For instance, the validation data may come from real-world user queries that cover unknown compositions of various domains. To remove the constraint on validation components, we assume that we can decompose the validation data into $K$ implicit domains whose losses are predictable with Eqn. 7, and we treat their proportions in the validation data $s_{1\dots K}$ as learnable parameters. This leads to the final form of our data mixing laws [^6] We note that the final forms of our data mixing law resemble a multilayer perception (see the computation graph Fig. 10). We include further discussion and implementation details in Appendix C.. With this perspective, we fit a data mixing law with the overall losses end to end.
@@
Fig. 5 shows our experiments where we train language models on the 5 coarse-grained domains of Pile [^7] The Pile contains 22 fine-grained domains which are collected into five coarse-grained domains, i.e., academic, internet, prose, dialogues, and misc, where misc include Github and the DeepMind Mathematics Dataset which are symbolic content. We do not experiment on fine-grained domains for their limited number of tokens available. and evaluate a validation set mixed with these 5 domains. We compare the errors obtained by implicit domain aggregation with different numbers of implicit domains to those obtained by explicit domain aggregation. We find that applying implicit domain aggregation and setting the number of implicit domains no smaller than the actual one (5 in the experimented case) results in lower errors than explicit domain aggregation. Moreover, the error remains low as we set the number of implicit domains much larger. This verifies the prediction accuracy of our implicit domain aggregation strategy for data mixing law and the number of implicit domains $K$ can be a large number without careful tuning [^8] We set $K=30$ if not stated in the rest of this paper..
@@ 4 Nested scaling laws predict losses trained on various mixtures using only small-scale experiments
1:Input: Validation data $D_{val}$, training data of $M$ domains $\{D_{m}\}_{m=1}^{M}$, target training steps $S_{target}$, target model size $N_{target}$, target mixture to predict $\bm{r}_{target}$, training steps to fit the step laws $S_{0}$, model sizes to fit the size laws $\{N_{i}\}$, and $N$ data mixtures $\{\bm{r}_{i}\}_{i=1}^{N}$ to fit.
@@
6:         Fit training step law to predict $L(N_{j},S_{target},\bm{r}_{i})$
7:     end for
8:     Fit model size scaling law to predict the loss $L(N_{target},S_{target},\bm{r}_{i})$
9:end for
10:Fit the data mixing law to predict $L(N_{target},S_{target},\bm{r}_{target})$
@@ 4.1 A Pipeline for Loss Predictions
While data mixing laws enable us to predict the performance of models trained on unseen mixtures, the requirement to fit the laws involves training multiple models across diverse mixtures with model sizes and token counts identical to the target ones. Furthermore, we must repeat this process for each target model size and training dataset [^9] An idea is to transfer the optimized training mixture on small models trained with few tokens to the training of large models and large volumes of training data. Nevertheless, in Appendix A, we show that the partial order of data mixture concerning model performance varies as the model size and number of trained tokens change. Therefore, the optimal mixture at experimented scales can be suboptimal at the target scale. . This results in expensive costs thus hindering the practical value of our data mixing laws.
We thus wonder whether we can obtain the losses of different mixture proportions without training at large scales. Fortunately, this idea gains endorsement from existing experiences that verify the impressive extrapolation of scaling laws of training steps and model sizes. In particular, OpenAI (2023) predicts the loss of the target model with merely 1,000×1,000$\times$1 , 000 ×10,000×10,000$\times$10 , 000 × less compute. As a consequence, we can train small models with few training steps on different mixtures, and fitting scaling laws on them to estimate the losses of the target model size and the target number of training steps. We can then use the predicted losses to fit a data mixing law and search for the optimal mixture. The proposed pipeline is illustrated in Fig. 1 with details depicted in Alg. 1.
We note that the nested use of scaling laws is still not a widely adopted strategy in existing literature to the best of our knowledge. Most studies try connecting newly found scaling laws to existing ones by incorporating variables into the latter (Hashimoto, 2021; Frantar et al., 2023; Aghajanyan et al., 2023). Although unifying of different scaling laws is attractive, the added complexity to the interaction among multiple variables makes identifying the function form extremely challenging and hard to verify. Pioneering efforts that derive scaling laws with multiple concerning factors (model sizes and the numbers of training data) (Kaplan et al., 2020; Hoffmann et al., 2022) make assumptions upon the theory of error decomposition (Harville, 1985; Domingos, 2000) but their outcomes are still not widely verified in large-scale practice. From this perspective, we consider that our nested use of scaling laws offers a new technical routine of using scaling laws, which circumvents the huge challenge of complicating existing laws, making the discovery and application of scaling laws more accessible.
@@ 4.2 Experiment
To reach low prediction errors with a limited number of experiment runs, we select the mixtures for experimentation by leveraging the fact that mixture proportion terms are represented as exponential functions within our data mixing law. Specifically, we enumerate candidate mixtures by double-diminishing the proportion for each training domain, starting from the maximum available one that does not use up all the domain tokens. In this way, the losses of each (implicit) validation domain are distributed evenly. We then sample 40 mixtures from all the candidates and train the smallest 70M models. We resample groups of 20 mixtures from them to fit the data mixing law and select the group that reaches minimum prediction errors on all 40 samples as our final set of mixtures to run our pipeline. For more details, please refer to Appendix B.2.
@@ Results.
Fig. 6 shows the default mixture of RedPajama used in Touvron et al. (2023a) and the optimized mixture obtained from Alg. 1 with their performance on the validation data. The loss predictions are in Fig. 7. We find that:
  • The pipeline optimizes performance effectively. The model trained on the optimized mixture can achieve a performance comparable to the one trained on the default mixture with only 73% steps. It eventually attains a performance that requires 48% more steps if trained using the default mixture. This indicates the effectiveness of our pipeline in mixture optimization.
  • Predictions provide references for the balance of model capabilities. Prediction results in Fig. 7 demonstrate plausible reference on the relative scale of losses on both the overall validation data and different validation domains. While the overall loss helps optimize the overall performance, losses on different domains show model capabilities in various aspects. We suppose that this facilitates further mixture curating for the nuanced balance of model capabilities.

### v2

@@ 1 Introduction
Pretraining data for large language models (LLMs) are typically a mixture of multiple domains, varying from English to minority languages (Doddapaneni et al., 2021; Li et al., 2023), from casual dialogs to formal academic writings (Taylor et al., 2022), and from texts to modalities like images and speeches (Zhan et al., 2024), among others. These data interplay with each other, showing complex relationships including facilitation, being unrelated, or conflict (Guo et al., 2024). This necessitates adjusting the mixture proportions of training data to balance the model capabilities while harnessing synergies across domains, thus enhancing the competence of the outcome models, as highlighted by extensive practices (Gururangan et al., 2020; Zhou et al., 2023; Xie et al., 2024a; Fan et al., 2024).
Nonetheless, it remains elusive to figure out an ideal training data mixture. Most existing practices tune the mixture through heuristics to upsample a proportion of high-quality or underrepresented data without disclosing the concrete criteria in detail (Gao et al., 2020; Touvron et al., 2023a; Bai et al., 2023; Bi et al., 2024). While some studies propose automatic algorithms to qualitatively optimize data mixture (Xie et al., 2024a; Fan et al., 2024), it is hard to predate the effect of these strategies before the actual training run. In contrast, encouraged by advances in scaling laws that show model losses on a given set of evaluation data are quantitatively predictable for a wide range of variables (Kaplan et al., 2020; Hoffmann et al., 2022), we wonder whether this also holds for mixture proportions, so that we can estimate the outcome model performance given any mixture before actually training on them, including the desired one that reaches minimum loss.
In this paper, we answer this proposition affirmatively. The intuition is that predicting the performance of unseen data mixture only involves interpolating among seen mixtures because the proportions are bounded between 0 and 1. For this reason, numerous function forms can lead to descent prediction accuracy, among which we try to find a simple one. In particular, we find that, given a mixture of $M$ domains, an exponential function over the linear combination of the proportions, i.e.,
@@ 2 Background
We briefly review the pretraining process of large language models and summarize key findings from neural scaling laws, then we formalize the problem we study. Further related works are in App. A.
@@
For a wide range of factors $x$, scaling laws (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022) show that their effect on the losses $L$ follows power laws
@@
where $c$, $k$, and $\alpha$ are parameters to fit and $x$ can be model sizes, numbers of training data, training steps, and the amount of computation. Previous experience (Alabdulmohsin et al., 2022; OpenAI, 2023; Bi et al., 2024; Su et al., 2024) highlights the impressive predictability of scaling laws. Specifically, Eqn. 3 fitted on a collection of small models, training data, or computation can extrapolate to precisely predict the test loss of larger cases over orders of magnitudes. This enables practitioners to estimate the performance of a pretrained large language model without actually finishing the expensive runs. Recent development further shows various functional relationships between the performance of language models and a broader range of factors, including transfer learning (Hernandez et al., 2021), sparse architectures (Frantar et al., 2023), and repeated data (Muennighoff et al., 2024), consolidating the predictability of language model performance.
@@ 3.1 Pilot Study on Domain Losses under Two-domain Mixtures
Results in Fig. 2 reveal the quantitative predictability of domain losses given the domain proportions. We encouragingly find that, for checkpoints with the same size and trained with the same number of steps, after subtracting a shared constant [^2] The constant term, known as irreducible loss, arises from finite training data and the entropy of the evaluation data theoretically (Bishop, 2006; Henighan et al., 2020)., their domain losses in the log scale demonstrate a linear relationship to the domain proportion. This holds for both domains in our experiments. The result indicates that with other factors fixed, the domain losses of a pretrained language model regarding the domain proportion precisely fit into an exponential law [^3] While power laws are more common in existing studies on scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022), we do not consider it for its ill-posed properties that the function value blows up when the variable, mixture proportion in our case, approaches 0. This contradicts the observations that the losses remain low (e.g., no more than 10) with the generalization between domains.
@@ 3.2 Extension to Domain Losses Trained on Multi-domain Mixtures
We train on the mixtures of GitHub, Pile-CC, and Books3 subset from the Pile for a total of 30B tokens and evaluate the model on the three domains, respectively. For specific mixtures, we grid search from $\{0,0.125,0.25,\dots,0.875,1\}^{3}$ and retain valid ones in which three proportions sum up to 1 and do not use up all tokens in any of the domains, which results in 32 mixtures in total.
@@
  • Symmetry. Any exchanging of variables should not change the functional form as we should not incorporate any domain-specific bias.
Together, the two principles lead to candidate functions that replicate the exponential term in Eqn. 6 for each training domain and combine them through operations that adhere to commutative law.
@@
Meanings of the coefficients.
To provide more intuition, we discuss the meanings of the coefficients in Eqn. 7. In general, $k_{i}>0$, thus the exponential term is always positive and the prediction loss is strictly greater than the constant $c$. Hereby, $c_{i}$ represents losses that are not reducible by adjusting the data mixture. $t_{ij}$, depending on both training domain $i$ and validation domain $j$, shows the interaction between them. A negative $t_{ij}$ indicates that training data of domain $j$ helps reduce validation loss on domain $i$ and vice versa.
Figure 4: The interaction between different training and validation domains on the Pile. Each boxes are fitted normalized $t_{ij}$ from Eqn. 7. We normalize the value by $t_{ij}$ with the maximum absolute value for each validation set $i$ (i.e., $t_{ij}/t_{i,\arg\max_{j}{|t_{ij}|}}$). A larger value (greener) indicates more mutual facilitation.
Patterns of the coefficients.
We visualize normalized $t_{ij}$ of training and validating on the mixture of 5 coarse-grained domains of the Pile.The relationship between domains can be categorized into 3 types. Being unrelated: The figure shows a highly sparse pattern where most of the domains have little relationship to each other and the validation performance of a domain is dominated by training data of the same domain, which supports the intuitive progressive mixture tuning strategy that adds data for underperforming capability during training (Team, 2023). Meanwhile, we also observe facilitation (e.g., training dialogue for the internet) and conflict (e.g., training symbolic data for prose) between domains, which indicates the room for leveraging domain interaction to enhance model performance.
[…]
3.3 Predicting Language Modeling Performance of Any Validation Mixture
@@
A remaining limitation is that we still need to acquire the components of validation data $s_{1\dots K}$ in advance. This can be inconvenient if the validation set is collected separately from the training ones. For instance, the validation data may come from real-world user queries that cover unknown compositions of various domains. To remove the constraint on validation components, we assume that we can decompose the validation data into $K$ implicit domains whose losses are predictable with Eqn. 7, and we treat their proportions in the validation data $s_{1\dots K}$ as learnable parameters, leading to the final form of our data mixing laws [^5] We note that the final forms of our data mixing law resemble a multilayer perception (see the computation graph Fig. 14). We include further discussion and implementation details in Appendix E.. With this perspective, we fit a data mixing law with the overall losses end to end.
Introducing implicit domains may draw concerns about the number of fitting samples exploding with the number of parameters to fit and questions on deciding the number of implicit domains without knowing the oracle. We study and discuss their impact, respectively.
Figure 6: The mean absolute validation errors of Eqn. 8 fitted with different numbers of samples for training mixtures containing different numbers of training domains. For each number, we resample and select the batch of fitting mixtures that reach lowest errors.
Do we need quadratic number of samples to fit Eqn. 8 as the number of domain grows? No. The parameters in Eq.8 scale as $\mathcal{O}(M\times K)$, where $M$ and $K$ represent training and implicit validation domains. Nevertheless, as shown in Fig.6, the quadratic growth in the number of parameters does not translate to quadratic growth in sample requirements. We attribute this to the high sparsity in the parameters as fig.4 reveals, which allows us to fit the equation with substantially fewer samples when using appropriate regularization. While using more samples decreases prediction errors, the number of samples that reach a similar precision level does not grow dramatically. This may pave the way for applying implicit domain aggregations for cases with more training domains. Although concluding the exact number of samples required can be challenging due to the differences among training data, we can tune the fitting mixtures on the smallest experimented models, which is cheap and works well in practice (see Sec. 4.2 and App. D.3).
@@
How to choose the number of implicit domains? Set it larger than the oracle one. Fig. 7 shows our experiments where we train language models on the 5 coarse-grained domains of Pile and evaluate a validation set mixed with these 5 domains. We compare the errors obtained by implicit domain aggregation with different numbers of implicit domains to those obtained by explicit domain aggregation. We find that applying implicit domain aggregation and setting the number of implicit domains no smaller than the actual one (5 in the experimented case) results in lower errors than explicit domain aggregation. Moreover, the error remains low as we set the number of implicit domains much larger. This verifies the prediction accuracy of our implicit domain aggregation strategy for data mixing law and the number of implicit domains $K$ can be a large number without careful tuning [^6] We set $K=30$ if not stated in later experiments..
@@ 4 Nested scaling laws predict losses trained on various mixtures using only small-scale experiments
1:Input: Validation data $D_{val}$, training data of $M$ domains $\{D_{m}\}_{m=1}^{M}$, target training steps $S_{target}$, target model size $N_{target}$, target mixture to predict $\bm{r}_{target}$, training steps to fit the step laws $S_{0}$, model sizes to fit the size laws $\{N_{j}\}_{j=1}^{K}$, and $N$ data mixtures $\{\bm{r}_{i}\}_{i=1}^{N}$ to fit.
@@
6:         Fit training step scaling law $L(S)$ with $L(N_{j},S<S_{0},\bm{r}_{i})$
7:         Predict $L(N_{j},S_{target},\bm{r}_{i})\leftarrow L(S=S_{target})$
8:     end for
9:     Fit model size scaling law $L(N)$ with $L(N_{1\dots K},S_{target},\bm{r}_{i})$
10:     Predict $L(N_{target},S_{target},\bm{r}_{i})\leftarrow L(N=N_{target})$
11:end for
12:Fit the data mixing law $L(\bm{r})$ with $L(N_{target},S_{target},\bm{r}_{1\dots N})$
13:predict $L(N_{target},S_{target},\bm{r}_{target})\leftarrow L(\bm{r}_{target}$
@@ 4.1 A Pipeline for Loss Predictions
While data mixing laws enable us to predict the performance of models trained on unseen mixtures, fitting the laws directly on target scales is unaffordably expensive. Firstly, fitting the laws involves training multiple models across diverse mixtures with model sizes and token counts identical to the target ones. Furthermore, we must repeat this process for each target model size and training dataset [^7] An idea is to transfer the optimized training mixture on small models trained with few tokens to the training of large models and large volumes of training data. Nevertheless, some recent studies highlight that the rankings of the data mixture vary as the model size and number of trained tokens change (Goyal et al., 2024; Kang et al., 2024; Covert et al., ). Therefore, the optimal mixture at experimented scales can be suboptimal at the target scale. . Such expensive costs hinder the practical value of our data mixing laws.
We thus wonder whether we can obtain the losses of different mixture proportions without training at large scales. Fortunately, this idea gains endorsement from existing experiences that verify the impressive extrapolation of scaling laws of training steps and model sizes. In particular, OpenAI (2023) predicts the loss of the target model with merely 1,000×1,000$\times$1 , 000 ×10,000×10,000$\times$10 , 000 × less compute. Consequently, we can train small models with few training steps on different mixtures, and fitting scaling laws on them to estimate the losses of the target model size and the target number of training steps. We can then use the predicted losses to fit a data mixing law and search for the optimal mixture.
We illustrate the proposed pipeline in Fig. 1 with details depicted in Alg. 1. Scaling laws in our pipeline are largely based on the function forms of Chinchilla Scaling Laws (Hoffmann et al., 2022), i.e., $L(N,D)=E+\frac{A}{N^{\alpha}}+\frac{B}{D^{\beta}}$, where $N$ is the model size and $D$ is the number of training data. Under fixed batch sizes, we can treat the number of training data as the number of training steps $S$ as well. Notably, we do not directly fit the complete Chinchilla Scaling Law with two variables as we find it practically unstable to fit such many parameters simultaneously in our preliminary study, similar to the findings in Besiroglu et al. (2024). Instead, we decompose the law into two power laws for training steps $L(S)=E_{1}+\frac{B}{S}$ and model sizes $L(N)=E_{2}+\frac{A}{N}$, respectively. We first fit power laws of training steps to predict model performance with more training data then fit power laws of model sizes to predict the performance when scaling up models. We empirically find this routine stable.888 See Appendix D.2 for our preliminary verification. We notice some recent efforts try to investigate democratizing the implementation of predictions with scaling laws to facilitate applications (Su et al., 2024; Porian et al., 2024). While we illustrate our pipeline with the nested use of scaling laws, other implementations of scaling law predictions are also feasible and complementary to our method. .
@@ 4.2 Experiment
To reach low prediction errors with a limited number of experiment runs, we select the mixtures for experimentation by leveraging the fact that mixture proportion terms are represented as exponential functions within our data mixing law. Specifically, we enumerate candidate mixtures by double-diminishing the proportion for each training domain, starting from the maximum available one that does not use up all the domain tokens. In this way, the losses of each (implicit) validation domain are distributed evenly. We then sample 40 mixtures from all the candidates and train the smallest 70M models. We resample groups of 20 mixtures from them to fit the data mixing law and select the group that reaches minimum prediction errors on all 40 samples as our final set of mixtures to run our pipeline. For more details, please refer to Appendix D.3.
@@ Results.
Our pipeline optimizes language modeling performance effectively. Fig. 8 shows
Figure 9: The language modeling performance of different data mixtures. All models are 1B models trained for 100B tokens with the same hyperparameters and validated on the validation set of the Pile. Specific proportions are in Fig. 21.
the default mixture of RedPajama (Touvron et al., 2023a) in and the optimized mixture obtained from Alg. 1 with their performance on the validation data. The model trained on the optimized mixture can achieve a performance comparable to the one trained on the default mixture with only 73% steps. It eventually attains a performance that requires 48% more steps if trained using the default mixture. This indicates the effectiveness of our pipeline in mixture optimization [^9] The loss predictions are in Fig. 22, which shows the predictions are plausibly accurate and are consistent with the rankings of actual runs..
We also compare our optimized data mixture to previous optimization algorithms, which provide qualitative optimization. Specifically, we compare our method to DoGE (Fan et al., 2024) and DoReMi (Xie et al., 2024a). For DoGE, we compare both their universal generalization setting which assumes i.i.d. between training and validation data, and the OOD setting which optimizes for a given validation set, similar to ours. For DoReMi, which only works for universal optimization that ignores the validation data, we experiment on both a mixture optimized exactly on RedPajama and a mixture adapted from the one optimized on the Pile using the domain overlap between RedPajama and the Pile. More specific details on obtaining these data mixtures are in Appendix F.4. As shown in Fig. 9, our method finds the mixture that reaches the lowest losses for the same model sizes trained with the same data budgets. This further verifies the effectiveness of our method.